OpenStudy (anonymous):

Solve this initial value problem... \y''-y=t-2 \space ; \space y(2)=3 \space , \space y'(2)=0\]

OpenStudy (anonymous):

\[y''-y=t-2 \space ; \space y(2)=3 \space , \space y'(2)=0\]

OpenStudy (anonymous):

Time shift... w(t)=y(t+2) w'(t)=y'(t+2) w''(t)=y''(t+2)

OpenStudy (anonymous):

Check me so far please

OpenStudy (anonymous):

@zepdrix

OpenStudy (anonymous):

\[w''-w=t \space ; \space w(0)=3 \space , \space w'(0)=0\]

OpenStudy (anonymous):

After partial fractions I end up with...\[\large -\frac{ 1 }{ s^2 }+3\frac{ s }{ s^2-1 }-\frac{ 1 }{ s^2-1 }\]

OpenStudy (anonymous):

inverse Laplacing I get\[-t+3\cos(-t)-\sin(-t)\]

OpenStudy (anonymous):

Back of the book says\[2-t+e^{2-t}+2e^{t-2}\]

OpenStudy (anonymous):

I am missing the final step looks like. I have to reverse the time shift

zepdrix (zepdrix):

Hmm I'm not familiar with the time shift thing.. (Just starting to learn this stuff past friday). Is it because we need our initial conditions to start at time 0?

OpenStudy (anonymous):

Ya. Initial conditions must be at zero for Laplace

OpenStudy (anonymous):

if w(t)=y(t+2), then y(t)=w(t-2). It looks like I might have made a mistake during the problem. I shouldn't have any trig in here. I'm going to rework it. I'll post the work if you're interested to see a time shift @zepdrix

zepdrix (zepdrix):

Mmm ok i think i see your mistake, sec lemme check my work a sec.

zepdrix (zepdrix):